OpenStudy (anonymous):

What is x to the second power - 5x + 6 = 0

OpenStudy (anonymous):

Do you mean x^2 - 5x + 6?

OpenStudy (anonymous):

You need to use the quadratic formula.

OpenStudy (anonymous):

Yea Ga

OpenStudy (anonymous):

\[x = \frac{ -b \sqrt{b^{2} - 4ac}}{ 2a }\]

OpenStudy (anonymous):

im not good in math !

OpenStudy (anonymous):

I'll help you. :)

OpenStudy (anonymous):

Okay, a quadratic equation looks like this: \[ax^{2}+bx+c=0\]

OpenStudy (anonymous):

So for your equation, \[x^{2} - 5x + 6 = 0\] \[a = 1, b = -5, c = 6\]

OpenStudy (anonymous):

So we plug that into the quadratic formula, and then solve. I'm going to need you to put in some input so I don't just give you the answer. Fair right?

OpenStudy (anonymous):

\[x = \frac{ 5 \sqrt{5^{2} - 4(1)(6)}}{ 2(1) }\]

OpenStudy (anonymous):

So, the first step is to simplify what's under the square root. \[\sqrt{5^{2} - 4(1)(6)}\]

OpenStudy (anonymous):

do you understand so far?

OpenStudy (anonymous):

@DestinyWay16

OpenStudy (anonymous):

I'll explain how to do this but I need to make sure you're there and that you understand what I'm doing.

OpenStudy (anonymous):

no i dnt

OpenStudy (anonymous):

Okay, what do you not understand so far: You don't know how I plugged in the variables into the equation? You don't know how to solve: \[\sqrt{5^{2} - 4(1)(6)}\]

OpenStudy (anonymous):

@DestinyWay16

OpenStudy (anonymous):

no

OpenStudy (anonymous):

no what?